# Dynamic In-context Learning with Conversational Models for Data Extraction and Materials Property Prediction

### v1

@@ Abstract
The advent of natural language processing and large language models (LLMs) has revolutionized the extraction of data from unstructured scholarly papers. However, ensuring data trustworthiness remains a significant challenge. In this paper, we introduce PropertyExtractor, an open-source tool that leverages advanced conversational LLMs like Google gemini-pro and OpenAI gpt-4, blends zero-shot with few-shot in-context learning, and employs engineered prompts for the dynamic refinement of structured information hierarchies – enabling autonomous, efficient, scalable, and accurate identification, extraction, and verification of material property data. Our tests on material data demonstrate precision and recall exceeding 93% with an error rate of approximately 10%, highlighting the effectiveness and versatility of the toolkit. We apply PropertyExtractor to generate a database of 2D material thicknesses, a critical parameter for device integration. The rapid evolution of the field has outpaced both experimental measurements and computational methods, creating a significant data gap. Our work addresses this gap and showcases the potential of PropertyExtractor as a reliable and efficient tool for the autonomous generation of diverse material property databases, advancing the field.
@@
The advent of generative artificial intelligence, particularly through advances in large language models (LLMs) and natural language processing (NLP), represents a transformative shift in the ability to harness unstructured data across diverse fields, such as material science. The abundance of data, which includes journal articles, patents, and theses, could be overwhelming for manual data extraction, especially due to the rapid pace of publication and its voluminous nature. Efficiently extracting and utilizing this information for material characterization and discovery remains a significant challenge. Traditional methods for data extraction from unstructured texts in materials science have relied on labor-intensive setups, including the development of parsing rules and the identification of specific phrases, necessitating extensive model fine-tuning and re-training. Such processes are both resource-consuming and time-intensive, often resulting in systems that are rigidly specialized for narrow tasks KONONOVA2021102155 ; Gilligan2023 . Recent advancements in the generative pretrained transformer series of LLMs, noted for their exceptional capabilities in generating coherent and contextually relevant text, summarizing content, and identifying relationships between concepts, offer a promising solution. These conversational LLMs automate the extraction of relevant information from extensive document collections, facilitating the creation of material property databases and accelerating the discovery of new materials Polak2024 ; D3DD00113J ; Dagdelen2024 .
@@
Recent advancements in the field of materials science have significantly benefited from the development of NLP techniques aimed at extracting valuable information from unstructured research papers. These efforts have laid the groundwork for sophisticated data analysis and extraction methods, allowing researchers to glean insights that were previously inaccessible Polak2024 ; D3DD00113J ; Gilligan2023 ; D3DD00019B . Despite these advances, challenges remain in achieving high fidelity data and context-aware data extraction, particularly when dealing with complex physical properties and diverse scientific terminologies. Huang & Cole fine-tuned a BERT model on battery-related publications to enhance a battery database, employing a “question and answer” (Q/A) strategy to extract specific device-level information. However, their method struggles with passages containing multiple device information and requires extensive pre-training on battery research papers before fine-tuning for the Q/A tasks doi:10.1021/acs.jcim.2c00035 . More recently, Zheng et al. developed a prompt-engineering approach, ChemPrompt with ChatGPT, focused on transforming text into tabular forms and summarizing scientific papers, leveraging the vast pretraining corpus to create semi-structured summaries doi:10.1021/jacs.3c05819 . Similarly, studies by Castro and Pimentel explored ChatGPT’s baseline chemistry knowledge and found that without advanced prompt engineering, the model’s performance was suboptimal on straightforward chemistry tasks doi:10.1021/acs.jcim.3c00285 . The work of Xie et al. expanded the use of LLMs in materials science, fine-tuning them on a broad corpus for diverse tasks such as Q/A, inverse design, classification, and regression Xie2023DARWIN . Despite these advancements, these approaches often fail to extract structured representations of complex hierarchical entity relationships and do not generalize beyond the limitations of the pretraining corpus. Recent works also showed that the performance of LLM-based architectures in extracting material property information can be improved via engineered prompts Dagdelen2024 ; Polak2024 . While these advances have significantly enhanced the structured information extraction of material property data, there is still an inherent issue of scalability and transferability, including challenges in extracting structured representations of complex hierarchical entity relationships generalizing outside of the pretraining corpus.
@@
In this paper, we introduce a blended dynamic zero-shot-few-shot in-context learning approach. Here, task-specific instructions, such as “Yes” and “No” responses (zero-shot learning) are combined with non-prescriptive guidance (few-shot learning) that dynamically incorporates accurately performed tasks into the model, enhancing and providing a closed feedback loop for both scalability and predictability. The concept of in-context learning (ICL) has emerged as a central paradigm for task adaptation in LLMs Dong2023InContextLearning , fundamentally enabling the model to adapt its behavior based on provided examples rather than undergoing resource-intensive fine-tuning of its internal parameters. ICL effectively leverages the “context” embedded within the model’s prompt to adapt the LLM to specific downstream tasks, spanning a spectrum from zero-shot learning (where no additional examples are provided — task descriptive instructions) to few-shot learning (where several examples are offered). Leveraging the inherent semantic similarity, our approach dynamically selects contextually relevant outcomes from a set of accurately prior predicted tasks, thereby enhancing prompt engineering and scalability for the predictability of the large language model-based toolkit. Such an approach is pivotal in materials science, particularly for extracting critical property data such as those of novel two-dimensional (2D) materials Novoselov2012 ; kastuar2022efficient ; NAJMAEI2020110 ; Khanmohammadi2024 , which generally lack generalizable property databases compared to conventional bulk materials.
@@
We implement this approach in the open-source code, PropertyExtractor, a Python-based computational framework designed to generate structured material property datapoint quadruplets of material, property value, original unit, method. It employs a sophisticated combination of engineered prompts and dynamic zero-shot-few-shot in-context learning to identify data-rich sentences, whether these appear within tables or as single or multiple data values within text. This approach not only extracts data but also verifies its accuracy, addressing common issues associated with factual inaccuracies typically observed in large language model responses. PropertyExtractor achieved significant improvements in model scalability, realizing a precision of $\sim 96\%$, a recall of $94\%$, an accuracy of 90%, an F1-score of 95%, and an error rate of $\sim$10on a constrained dataset of the thickness of 2D materials. These performance metrics surpass those reported in earlier studies Polak2024 ; D3DD00113J ; Gilligan2023 , highlighting its potential utility in both academic and industrial settings. Moreover, PropertyExtractor is designed as a software and leverages online LLM APIs, which allow users to train bespoke models without extensive knowledge of the internal workings of LLMs or Python skill. The user may simply treat the PropertyExtractor architecture as a black-box that transforms passages into precisely-formatted, structured summaries of material property data, thus enabling researchers with little NLP experience to use the toolkit. While PropertyExtractor is currently implemented for OpenAI gpt and Google gemini, its architecture is deliberately designed for easy adaptation to incorporate any conversational LLM. This flexibility is achieved by modularizing the component that interfaces with the LLM, allowing for simple modifications of a specific subroutine to accommodate alternative models. Such an adaptable system ensures that PropertyExtractor not only keeps pace with the rapid advancements in the field of LLMs but also leverages these improvements to enhance its functionality. The evolving landscape of LLMs development suggests a future where the continual improvements in model performance could further amplify the efficacy of PropertyExtractor. Drawing parallels from the field of image generation, where prompt engineering has become a standard practice to secure high-quality results, a similar trend is expected in data extraction. Our approach, with its foundation in dynamical in-context learning, prompt engineering, and conversational prompts, positions PropertyExtractor as a versatile tool adaptable to the forthcoming generations of LLMs, ensuring high-quality data extraction across diverse applications.
I Results and Discussions
I.1 Data Extraction Workflow
@@
Figure 1 shows a concise workflow of the integrative unstructured data collection via API and structured data extraction with the PropertyExtractor. In Figure 2, we present a detailed workflow illustrating the essential processes and steps for extracting material properties from unstructured research papers. The data extraction process begins with data preparation. This stage involves programmatically gathering relevant scientific literature using various APIs and refining the content to create clean, unstructured data suitable for the conversational language model. The data curation is specifically designed to target texts mentioning the principal material property and may include an optional set of keywords related to the specific property, using an extensive set of keywords to cover the diverse terminology within the field. While the keywords are optional, they are essential for ensuring that the extracted data is relevant to the specified material property. For example, extracting energy band gap specific to 2D materials could have “band gap” as the property and “2D material” as one of the keywords. The initial process of the data extraction involves two distinct stages. The first stage is the data collection phase, where scientific papers are programmatically gathered using API. This is followed by the data preparation phase, where the collected papers are cleaned to remove XML/HTML tags and other syntactical clutter and then segmented into individual sentences for further processing. Following the cleaning of unstructured data, the next crucial step involves incorporating the LLM-based model, which is enhanced by dynamic zero-shot-few-shot in-context learning. This advancement refines the LLM’s conversational interactions with the data using contextually rich prompts. We found that integrating zero-shot and few-shot in-context learning with dynamic refinement through feedback on recently, accurately predicted outcomes enables the LLM to adaptively refine the data extraction process. This strategy significantly enhances extraction accuracy by minimizing errors and preventing data hallucinations. It effectively addresses the challenges posed by the diverse structures, complexity, and terminology found in unstructured scientific literature, ensuring meticulously calibrated responses for high-fidelity data extraction. The critical steps of this process are broken down as follows:
@@
    • Dynamically adjusting prompts. As the model processes the data, it dynamically adjusts the prompts based on recent analyzes. This adaptation captures the complexity and variability inherent in scientific texts, tailoring interactions to improve understanding and response accuracy.
    • Handling data variability. The model is equipped with strategies to efficiently manage both single- and multi-valued data points, enhancing its adaptability across various data extraction scenarios. For multi-valued scenarios, the model uses regex patterns and dynamic zero-shot-few-shot in-context learning to address the complexity of extracting and correctly associating each value with its respective material and property. This method begins with detecting the presence of multiple data points within a sentence, followed by tailored prompts that guide the model in differentiating and processing these values accurately. Additionally, the model employs prompts that encourage critical evaluation of its responses. This not only reduces errors but also significantly enhances the trustworthiness of the extracted data by ensuring that each data point is accurately captured and attributed.
@@
I.2 Performance Evaluation
We evaluated the performance of PropertyExtractor by generating a database for the thickness of 2D materials. This focus on thickness is driven by two main factors. First, despite its crucial importance for integrating 2D materials into devices, there is currently no database or computational framework for determining the thickness of these materials. The thickness of 2D materials significantly influences their electronic structure, optical properties, mechanical behavior, and chemical reactivity. For example, MoS2 exhibits a transition from an indirect to a direct bandgap when reduced to a single layer, enhancing its suitability for optoelectronic devices. Similarly, the mechanical properties of graphene and the surface-to-volume ratio, critical for catalysis and sensor applications, are thickness-dependent. This property is also vital for the performance of energy storage devices, impacting ion transport and storage capacity.
Second, the thickness dataset poses a significant challenge due to the potential for confusion with similar properties, such as layer count or interlayer spacing, which require precise differentiation. Hence, it provides a robust testbed to evaluate the scalability and capability of PropertyExtractor, demonstrating its effectiveness in accurately discerning and extracting complex data amidst closely related variables.
@@ I.2 Performance Evaluation
Fulltexts or abstracts were processed using the Google gemini-pro API conversational language model within Python 3.10.12, employing our dynamic zero-shot-few-shot in-context learning and prompt engineering approach. This process yielded 1,221 raw data points, which, after cleaning, resulted in 1,015 data points. Further post-processing standardized these to 584 data points with units specified in Å. Among these, some records pertained to materials that might not be fully atomically thin 2D materials, despite initial keyword searches designed to exclude such instances. The inclusion of non-2D materials does not reflect a deficiency in the capabilities of our conversational LLM, as these data were present in the literature sourced. Restricting the database exclusively to atomically thin 2D materials yielded a final standardized database containing thickness values for about 43% of the dataset. The extracted thickness dataset is dominated by the well-known 2D materials such as graphene, transition metal dichalcogenides, BN, phospherene, etc.(Figure 3). Within this dataset, 68 thickness values corresponded to unique materials. It is important to note that in cases where texts explicitly mention 2D material thickness, the model extracted these instances with nearly 100% accuracy, demonstrating the robustness of PropertyExtractor in handling texts with less ambiguous keyword descriptions.
@@
Evaluation Methodology for PropertyExtractor. Given the lack of a ground truth database for the thickness of 2D materials, we developed a unique approach to evaluate the performance of PropertyExtractor. We curated a dataset from the scientific literature, extracting data to establish a ground truth. From approximately 12,000 fulltexts or abstracts, we selected 100 relevant ones containing thickness data. This selection process was targeted to ensure that the model only processes potentially useful data, optimizing computational resources and focusing on relevant texts. We defined our ground truth as 50 manually extracted data points from these texts, focusing on triplets comprising the material [chemical formula], its thickness value, and the measurement unit. This methodology allows for a precise evaluation by testing the model against a well-defined set of data, which is crucial for assessing the tool’s capability to accurately extract specific material properties from complex scientific texts. The validation process is carried out using both gemini-pro and gpt-4.
@@
Utilizing PropertyExtractor with gemini-pro, the metrics achieved are precision of 95.74%, recall of 93.75%, F1-score of 94.73%, accuracy of 90.00%, and an error rate of approximately 10.00%. These values indicate that PropertyExtractor correctly predicts a triplet 95.56% of the time (precision) and successfully identifies 93.48% of all relevant triplets (recall). The accuracy of 89.58% reflects the overall correctness of the tool, accounting for both positives and negatives. An F1-score of approximately 94.51% suggests a balanced model in terms of precision and recall. The error rate, at around 10.42%, demonstrates the model’s effectiveness, with a small proportion of the predictions being incorrect. Furthermore, when evaluations are specifically based on the material and standardized thickness values, the performance notably improves, exemplified by a precision of approximately 98%. We note that the accuracy using gpt-4 is similar to that obtained with gemini-pro, but gpt-4 performs better in the effective determination of the method used and also in discerning data that are provided as a range.
II Conclusions

### v2

@@ Abstract
The advent of natural language processing and large language models (LLMs) has revolutionized the extraction of data from unstructured scholarly papers. However, ensuring data trustworthiness remains a significant challenge. In this paper, we introduce PropertyExtractor, an open-source tool that leverages advanced conversational LLMs like Google gemini-pro and OpenAI gpt-4, blends zero-shot with few-shot in-context learning, and employs engineered prompts for the dynamic refinement of structured information hierarchies – enabling autonomous, efficient, scalable, and accurate identification, extraction, and verification of material property data. Our tests on material data demonstrate precision and recall that exceed 95% with an error rate of approximately 9%, highlighting the effectiveness and versatility of the toolkit. Finally, databases for 2D material thicknesses, a critical parameter for device integration, and energy bandgap values are developed using PropertyExtractor. Specifically for the thickness database, the rapid evolution of the field has outpaced both experimental measurements and computational methods, creating a significant data gap. Our work addresses this gap and showcases the potential of PropertyExtractor as a reliable and efficient tool for the autonomous generation of various material property databases, advancing the field.
@@
I Significance Statement
High-fidelity material property databases are essential for addressing various science and engineering challenges, including algorithm development and downstream tasks. Recent advancements in generative AI technology have significantly improved material property extraction from unstructured literature, but maintaining data accuracy and reliability remains a challenge. Here, we introduce a computational framework that addresses this issue by leveraging advanced large language models (LLMs), integrating zero-shot and few-shot in-context learning techniques for dynamic information hierarchy refinement. This facilitates autonomous, efficient, and scalable extraction and verification of material properties. The user-friendly architecture allows researchers to generate accurate databases with minimal LLM or programming expertise, thereby democratizing access.
II Introduction
@@
The advent of generative artificial intelligence, particularly through advances in large language models (LLMs) and natural language processing (NLP), represents a transformative shift in the ability to harness unstructured data across diverse fields, such as material science. The abundance of data, including journal articles, patents, and theses, could be overwhelming for manual data extraction, especially due to the rapid pace of publication and its voluminous nature. Efficiently extracting and utilizing this information for material characterization and discovery remains a significant challenge. Traditional methods for data extraction from unstructured texts in materials science have relied on labor-intensive setups, including the development of parsing rules and the identification of specific phrases, necessitating extensive model fine-tuning and re-training. Such processes are resource-consuming and time-intensive, often resulting in systems that are rigidly specialized for narrow tasks KONONOVA2021102155 ; Gilligan2023 . Recent advancements in the generative pretrained transformer series of LLMs, noted for their exceptional capabilities in generating coherent and contextually relevant text, summarizing content, and identifying relationships between concepts, offer a promising solution. These conversational LLMs automate the extraction of relevant information from extensive document collections, facilitating the creation of material property databases and accelerating the discovery of new materials Polak2024 ; D3DD00113J ; Dagdelen2024 .
@@
Recent advances in the field of materials science have significantly benefited from the development of NLP techniques that aim to extract valuable information from unstructured research papers. These efforts have laid the foundation for sophisticated data analysis and extraction methods, allowing researchers to gain insights that were previously inaccessible Polak2024 ; D3DD00113J ; Gilligan2023 ; D3DD00019B . Despite these advances, challenges remain in achieving high fidelity data and context-aware data extraction, particularly when dealing with complex physical properties and diverse scientific terminologies. Huang & Cole fine-tuned a BERT model on battery-related publications to enhance a battery database, employing a “question and answer” (Q/A) strategy to extract specific device-level information. However, their method struggles with passages containing multiple device information and requires extensive pre-training on battery research papers before fine-tuning for the Q/A tasks doi:10.1021/acs.jcim.2c00035 . More recently, Zheng et al. developed a prompt-engineering approach, ChemPrompt with ChatGPT, focused on transforming text into tabular forms and summarizing scientific papers, leveraging the vast pretraining corpus to create semi-structured summaries doi:10.1021/jacs.3c05819 . Similarly, studies by Castro and Pimentel explored ChatGPT’s baseline chemistry knowledge and found that without advanced prompt engineering, the model’s performance was suboptimal on straightforward chemistry tasks doi:10.1021/acs.jcim.3c00285 . The work of Xie et al. expanded the use of LLMs in materials science, fine-tuning them on a broad corpus for diverse tasks such as Q/A, inverse design, classification, and regression Xie2023DARWIN . Despite these advancements, these approaches often fail to extract structured representations of complex hierarchical entity relationships and do not generalize beyond the limitations of the pretraining corpus. Recent works also showed that the performance of LLM-based architectures in extracting material property information can be improved via engineered prompts Dagdelen2024 ; Polak2024 . While these advances have significantly enhanced the structured information extraction of material property data, there is still an inherent issue of scalability and transferability, including challenges in extracting structured representations of complex hierarchical entity relationships generalizing outside of the pretraining corpus.
@@
In this paper, we introduce a blended dynamic zero-shot-few-shot in-context learning approach. Here, task-specific instructions, such as “Yes” and “No” responses (zero-shot learning) are combined with non-prescriptive guidance (few-shot learning) that dynamically incorporates accurately performed tasks into the model, enhancing and providing a closed feedback loop for both scalability and predictability. In the zero-shot learning method, the model leverages its pre-existing (or “general”) knowledge and understanding to generate responses or outputs relevant to tasks on which it was not specifically trained, based solely on the instructions given in the prompt. For instance, determining whether the sentence “Graphene is a semiconductor” is an instance of misconception could involve the following prompt:
Prompt:
Classify whether or not the text below
expresses graphene feature skepticism:
Text: "Graphene is a semimetal, not a
semiconductor, because it lacks a bandgap."
Classification:
Recent works have shown that zero-shot learning applications using LLMs can yield reasonable results Polak2024 . Nevertheless, zero-shot learning relies heavily on the model’s “general” knowledge and may not perform optimally on specific downstream and domain-specific tasks, such as material property analysis and material science. Few-shot learning addresses these limitations by providing additional domain-specific examples to enhance the LLM’s understanding. The model then generalizes from these examples to perform the task effectively, even with minimal training data. Consequently, combining zero-shot and few-shot learning opens up new possibilities for employing LLMs in a broad range of applications without the need for extensive fine-tuning on specific tasks, making them more versatile, adaptable to various contexts inherent in material science, and scalable.
@@
The concept of in-context learning (ICL) has emerged as a central paradigm for task adaptation in LLMs Dong2023InContextLearning , fundamentally enabling the model to adapt its behavior based on provided examples rather than undergoing resource-intensive fine-tuning of its internal parameters. ICL effectively leverages the “context” embedded within the model’s prompt to adapt the LLM to specific downstream tasks, spanning a spectrum from zero-shot learning (where no additional examples are provided — task descriptive instructions) to few-shot learning (where several examples are offered). Leveraging the inherent semantic similarity, our approach dynamically selects contextually relevant outcomes from a set of accurately prior predicted tasks, thereby enhancing prompt engineering and scalability for the predictability of the large language model-based toolkit. Such an approach is pivotal in materials science, particularly for extracting critical property data such as those of novel two-dimensional (2D) materials Novoselov2012 ; kastuar2022efficient ; NAJMAEI2020110 ; Khanmohammadi2024 , which generally lack generalizable property databases compared to conventional bulk materials.
@@
We implement this approach in the open-source code, PropertyExtractor, a Python-based computational framework designed to generate structured material property datapoint quadruplets of material, property value, original unit, method. It employs a sophisticated combination of engineered prompts and dynamic zero-shot-few-shot in-context learning to identify data-rich sentences, whether these appear within tables or as single or multiple data values within the text. This approach not only extracts data but also verifies its accuracy, addressing common issues associated with factual inaccuracies typically observed in large language model responses. PropertyExtractor achieved significant improvements in model scalability, realizing a precision of approximately 96%, a recall of 94%, an accuracy of 90%, an F1-score of 95%, and an error rate of approximately 10% on a constrained dataset of the thickness of 2D materials. For the extraction of energy bandgap values, PropertyExtractor demonstrated even better performance metrics. It achieved a precision of 96.81%, a recall of 94.72%, an F1-score of 95.21%, an accuracy of 92.05%, and an error rate of approximately 7.95%. These results highlight the effectiveness and reliability of PropertyExtractor in accurately extracting critical material properties from scientific literature. These performance metrics surpass those reported in earlier studies Polak2024 ; D3DD00113J ; Gilligan2023 , highlighting its potential utility in both academic and industrial settings. Moreover, PropertyExtractor is designed as a software and leverages online LLM APIs, which allow users to train bespoke models without extensive knowledge of the internal workings of LLMs or Python skill. The user may simply treat the PropertyExtractor architecture as a black-box that transforms passages into precisely-formatted, structured summaries of material property data, thus enabling researchers with little NLP experience to use the toolkit. While PropertyExtractor is currently implemented for OpenAI gpt and Google gemini, its architecture is deliberately designed for easy adaptation to incorporate any conversational LLM. This flexibility is achieved by modularizing the component that interfaces with the LLM, allowing for simple modifications of a specific subroutine to accommodate alternative models. Such an adaptable system ensures that PropertyExtractor not only keeps pace with the rapid advancements in the field of LLMs but also leverages these improvements to enhance its functionality. The evolving landscape of LLMs development suggests a future where continual improvements in model performance could further amplify the efficacy of PropertyExtractor. Drawing parallels from the field of image generation, where prompt engineering has become a standard practice to secure high-quality results, a similar trend is expected in data extraction. Our approach, with its foundation in dynamical in-context learning, prompt engineering, and conversational prompts, positions PropertyExtractor as a versatile tool adaptable to the forthcoming generations of LLMs, ensuring high-quality data extraction across diverse applications.
III Results and Discussions
III.1 Data Extraction Workflow
@@
Figure 1 shows a concise workflow of the integrative unstructured data collection via API and structured data extraction with the PropertyExtractor. In Figure 2, we present a detailed workflow illustrating the essential processes and steps for extracting material properties from unstructured research papers. The data extraction process begins with data preparation. This stage involves programmatically gathering relevant scientific literature using various APIs and refining the content to create clean, unstructured data suitable for the conversational language model. The data curation is specifically designed to target texts mentioning the principal material property and may include an optional set of keywords related to the specific property, using an extensive set of keywords to cover the diverse terminology within the field. While the keywords are optional, they are essential for ensuring that the extracted data is relevant to the specified material property. For example, extracting energy bandgap specific to 2D materials could have “band gap” as the property and “2D material” as one of the keywords. The initial process of data extraction involves two distinct stages. The first stage is the data collection phase, where scientific papers are programmatically gathered using API. This is followed by the data preparation phase, where the collected papers are cleaned to remove XML/HTML tags and other syntactical clutter and then segmented into individual sentences for further processing. Following the cleaning of unstructured data, the next crucial step involves incorporating the LLM-based model, which is enhanced by dynamic zero-shot-few-shot in-context learning. This advancement refines the LLM’s conversational interactions with the data using contextually rich prompts. We found that integrating zero-shot and few-shot in-context learning with dynamic refinement through feedback on recently, accurately predicted outcomes enables the LLM to adaptively refine the data extraction process. This strategy significantly enhances extraction accuracy by minimizing errors and preventing data hallucinations. It effectively addresses the challenges posed by the diverse structures, complexity, and terminology found in unstructured scientific literature, ensuring meticulously calibrated responses for high-fidelity data extraction. The critical steps of this process are broken down as follows:
@@
    • Dynamically adjusting prompts. As the model processes the data, it dynamically adjusts the prompts based on recent analyses. This adaptation captures the complexity and variability inherent in scientific texts, tailoring interactions to improve understanding and response accuracy.
    • Handling data variability. The model is equipped with strategies to efficiently manage single- and multi-valued data points, enhancing its adaptability across various data extraction scenarios. For multi-valued scenarios, the model uses regex patterns and dynamic zero-shot-few-shot in-context learning to address the complexity of extracting and correctly associating each value with its respective material and property. This method begins with detecting the presence of multiple data points within a sentence, followed by tailored prompts that guide the model in differentiating and processing these values accurately. Additionally, the model employs prompts that encourage critical evaluation of its responses. This not only reduces errors but also significantly enhances the trustworthiness of the extracted data by ensuring that each data point is accurately captured and attributed.
@@
III.2 Performance Evaluation
In this study, we provide two materials property databases: the thickness of 2D-based materials and the energy bandgap of materials. The database for thickness has been post-processed to remove entries that do not belong to 2D materials. Accounting for the exact thickness, such as being a single-layer, is not feasible since this parameter is not reported in most literature. For the energy bandgap, we provide the raw output obtained using the PropertyExtractor. We evaluated the performance of PropertyExtractor by developing databases for the energy bandgap of materials and the thickness of 2D-based materials. The choice of these databases was intentional to address two critical aspects: testing the scalability of PropertyExtractor in obtaining hard-to-get material property features like the thickness of 2D-based materials and validating its performance on a well-established material property, the energy bandgap, which is readily available in many databases.
The focus on thickness is driven by two main factors. Despite its crucial importance for integrating 2D materials into devices, there is currently no database or computational framework for determining the thickness of these materials. The thickness of 2D materials significantly influences their electronic structure, optical properties, mechanical behavior, and chemical reactivity. For example, MoS2 exhibits a transition from an indirect to a direct bandgap when reduced to a single layer, enhancing its suitability for optoelectronic devices. Similarly, the mechanical properties of graphene and its surface-to-volume ratio, which are critical for catalysis and sensor applications, are thickness-dependent. This property is also vital for the performance of energy storage devices, impacting ion transport and storage capacity. Additionally, the thickness dataset poses a significant challenge due to the potential for confusion with similar properties, such as layer count or interlayer spacing, which require precise differentiation. Hence, it provides a robust testbed to evaluate the scalability and capability of PropertyExtractor, demonstrating its effectiveness in accurately discerning and extracting complex data amidst closely related variables.
@@ III.2 Performance Evaluation
Fulltexts or abstracts were processed using the Google gemini-pro API conversational language model within Python 3.10.12, employing our dynamic zero-shot-few-shot in-context learning and prompt engineering approach as implemented in PropertyExtractor. This process yielded 1,221 raw data points, which, after cleaning, resulted in 1,015 data points. Further post-processing standardized these to 584 data points with units specified in Å. Among these, some records pertained to materials that might not be fully atomically thin 2D materials, despite initial keyword searches designed to exclude such instances. The inclusion of non-2D materials does not reflect a deficiency in the capabilities of our conversational LLM, as these data were present in the literature sourced. Restricting the database exclusively to atomically thin 2D materials yielded a final standardized database containing thickness values for about 43% of the dataset. The extracted thickness dataset is dominated by the well-known 2D materials such as graphene, transition metal dichalcogenides, BN, phospherene, etc.(Figure 3). Within this dataset, 68 thickness values corresponded to unique materials. It is important to note that in cases where texts explicitly mention 2D material thickness, the model extracted these instances with nearly 100% accuracy, demonstrating the robustness of PropertyExtractor in handling texts with less ambiguous keyword descriptions.
To develop the energy bandgap database using PropertyExtractor, we applied a similar procedure as we did for the thickness of 2D-based materials. The only difference was that the primary and secondary keywords were replaced with “bandgap” and “band gap,” respectively. To obtain the source research articles, we performed a search query exclusively employed the PubMed API, which initially provided us with 12,102 independent entries. After processing the unstructured data from the PubMed API, we refined these entries down to 9,987 unique full texts or abstracts following further pre-processing steps. Utilizing the PropertyExtractor in conjunction with the Google Gemini-pro API conversational language model within Python 3.10.12, we then processed the 9,987 unique entries. This comprehensive processing resulted in 1,238 entries containing energy bandgap values and identified 561 unique materials.
@@
Evaluation Methodology for PropertyExtractor. Given the lack of a ground truth database for the thickness of 2D materials, we developed a unique approach to evaluate the performance of PropertyExtractor. We curated a dataset from the scientific literature, extracting data to establish a ground truth. From approximately 12,000 fulltexts or abstracts, we selected 100 relevant ones containing thickness data. This selection process was targeted to ensure that the model only processes potentially useful data, optimizing computational resources and focusing on relevant texts. We defined our ground truth as 50 manually extracted data points from these texts, focusing on triplets comprising the material [chemical formula], its thickness value, and the measurement unit. This methodology allows for a precise evaluation by testing the model against a well-defined set of data, which is crucial for assessing the tool’s capability to accurately extract specific material properties from complex scientific texts. Similar approach is used to establish the ground truth data specific to the energy bandgap. The validation process is carried out using both gemini-pro and gpt-4.
@@
We first evaluated the performance of PropertyExtractor in developing a database for the thickness of 2D materials. Utilizing PropertyExtractor with gemini-pro, the achieved metrics are as follows: precision of 95.74%, recall of 93.75%, F1-score of 94.73%, accuracy of 90.00%, and an error rate of approximately 10.00%. These values indicate that PropertyExtractor correctly predicts a triplet 95.56% of the time (precision) and successfully identifies 93.48% of all relevant triplets (recall). The accuracy of 89.58% reflects the overall correctness of the tool, accounting for both positives and negatives. An F1-score of approximately 94.51% suggests a balanced model in terms of precision and recall. The error rate, at around 10.42%, demonstrates the model’s effectiveness, with a small proportion of the predictions being incorrect. Furthermore, when evaluations are specifically based on the material and standardized thickness values, the performance notably improves, exemplified by a precision of approximately 98%. We note that the accuracy using GPT-4 is similar to that obtained with gemini-pro, but GPT-4 performs better in the effective determination of the method used and in discerning data that are provided as a range.
The performance metrics for extracting the energy bandgap values are slightly improved compared to the thickness data, which we attribute mainly to the well-defined nature of energy bandgap in scientific articles. The achieved metrics are precision of 96.81%, recall of 94.72%, F1-score of 95.21%, accuracy of 92.05%, and an error rate of approximately 7.95%. While a large automatically extracted database of bandgap values for materials has been developed previously,Dong2022 ; Kim2020 direct quantitative comparison is not straightforward. However, the histogram of values obtained from the previous database exhibits a very similar shape in terms of material types to the data obtained here, further supporting the robustness of our data. Moreover, many well-known semiconductors and their bandgap values are adequately captured in our database. It is important to note that the database of energy bandgap developed here is more diverse and heterogeneous, having been extracted from various first-principles calculations and diverse experimental techniques. This diversity in data sources underlines the robustness of PropertyExtractor in handling a wide range of information and providing reliable outputs.
IV Conclusions